A nationwide survey found out that the average time that college students spent on their personal computer is 10.5 hours per week. A random sample of 28 college students showed that they spent 8.5 hours per week using their computers with a standard deviation of 1.2 hours. Test whether the average number of hours spent by the 28 college students is significantly lower than the national average of 10.5 hours. Use a level of significance of 5%

1
Expert's answer
2022-06-10T11:48:27-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge10.5"

"H_1:\mu<10.5"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=27" and the critical value for a left-tailed test is "t_c =-1.703288."

The rejection region for this left-tailed test is "R = \{t:t<-1.703288\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{8.5-10.5}{1.2/\sqrt{28}}=-8.8192"


Since it is observed that "t=-8.8192<-1.703288=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=27" degrees of freedom, "t=-2.7106" is "p=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 10.5, at the "\alpha = 0.05" significance level.